OpenStudy (anonymous):

the air temperature decreases 12 degrees for every 4,000 feet increase in altitude.If a hiker climbs 3,000 feet, by how much can she expect the temperature to decrease?

OpenStudy (anonymous):

12/4000=x/3000

OpenStudy (anonymous):

can this be drawn lol im more of a visual person

OpenStudy (anonymous):

What do you mean? Draw what i just typed?

OpenStudy (anonymous):

no sorry wrong question

OpenStudy (anonymous):

im still confused though

OpenStudy (anonymous):

Cross multiply the equation 12/4000=x/3000 4000x=36000

OpenStudy (anonymous):

ohhhh thanks

OpenStudy (anonymous):

x=?

OpenStudy (anonymous):